OpenStudy (anonymous):

Find all triples (p, q, r) of primes such that p^3 + q^3 + r^3-3pqr divides 2016.

OpenStudy (anonymous):

I also knows there is exactly 36 factors of 2016.

ganeshie8 (ganeshie8):

Hint : \[p^3 + q^3 + r^3-3pqr = (p+q+r)(p^2+q^2+r^2-pq-qr-rp)\]

OpenStudy (anonymous):

I have tried bruteforcing, but I do not think that it is effective.

ganeshie8 (ganeshie8):

yeah looks tricky that hint seems useless..

OpenStudy (kainui):

\[p^3 + q^3 + r^3-3pqr = pqr(p^2+q^2+r^2-3)\] So p,q, and r must divide 2016 for starters, that might eliminate most possibilities

OpenStudy (kainui):

whoops

OpenStudy (kainui):

ok, pqr divides it regardless, by factorization is of course hilariously wrong

OpenStudy (kainui):

no wait I'm going insane this is all nonsense hahaha

OpenStudy (anonymous):

It couldn't be 1, since p,q,r must be prime.

ganeshie8 (ganeshie8):

\(p+q+r\) must be a divisor of \(3^2*7\) \[p+q+r \in \{1, 3,7, 9, 21, 63\}\]

ganeshie8 (ganeshie8):

right, eliminate 1

OpenStudy (anonymous):

and It can't be 3 either

ganeshie8 (ganeshie8):

Nice, all other combinations must be possible.. . recalling goldbach conjecture

OpenStudy (anonymous):

Also 2 is a factor of 2016, so shouldn't p+q+r be a divisor of 3^2*7*2^5, making p+q+r∈{,7,9,21,63} still incomplete?

ganeshie8 (ganeshie8):

in above case, we're assuming \(p+q+r\) is odd

ganeshie8 (ganeshie8):

that is true if none of \(p, q, r\) is \(2\)

ganeshie8 (ganeshie8):

what we're doing is more or less bruteforce only lets think of some other ways maybe..

OpenStudy (isaiah.feynman):

This looks like a Diophantine problem.

OpenStudy (anonymous):

A Diophantine is used if the answer is an integer, but in this case we need a prime number.

OpenStudy (isaiah.feynman):

A prime is an integer. :P

OpenStudy (anonymous):

But not all integers are prime

OpenStudy (isaiah.feynman):

The converse of my statement, which is true.

OpenStudy (kainui):

The primes can't all be the same, because then the expression is 0. What if two of the primes are the same? Then with \(p=r\) we have: \[(2p+q)(p-q)^2\] Since \(2016 = 2^53^27\) this restricts us to \(p-q\) or \(q-p\) being equal to 2 or 3 only. In other words, p and q are either twin primes OR 2 and 5. Of course, this is only for the cases where we have repeated primes.

OpenStudy (anonymous):

This could work if we could prove that there is no case in which all different primes are p,q, and r.

OpenStudy (kainui):

Yeah totally, that's what I'm sorta drifting towards. I just figured I'd try to prove some degenerate cases to get a feel for what kinds of stuff is going on. On an unrelated note, if we force the inequality \(p<q<r\) can that get us anything

OpenStudy (anonymous):

Are you attempting to use p<q<r to prove that at least two of the letters have to be same?

OpenStudy (kainui):

No, but that would be convenient. I was attempting to do it because it looks like the other cases where you have repeated values are pretty much handled, so we can focus solely on when the primes are all distinct

OpenStudy (anonymous):

Im pretty sure, the primes can't be distinct.

OpenStudy (kainui):

Well in that case if you are sure enough we can take this approach forward when you have one prime that occurs twice. Cause all twin primes greater than 3,5 are of the form of the littler one \(6n-1\) and the larger one \(6n+1\) so that's pretty restrictive I guess. Might be pretty straightforward to prove that this has no solutions or something idk

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

2,3,7 disproves that

OpenStudy (anonymous):

The good part is the fact that I could easily brute force for primes that are all distinct.

OpenStudy (kainui):

Yeah if you really wanted to solve it and not look for tricks, we could just write a program and be done with it haha. Then we could look for patterns from within the answer to retroactively discover solutions too lol.

OpenStudy (anonymous):

True.

OpenStudy (anonymous):

But because this is a math olympiad question, I know there is a trick.

OpenStudy (kainui):

Hmm. Since 2016 has 8 prime factors, that means when we factor that thing the total number of factors they could have gets split between the two. For instance if we can show that \(p^2+q^2+r^2-pq-pr-qr \ne 1\) then that limits the other thing, \(p+q+r\) to having at most 7 prime factors in it. Not terribly useful probably though haha

OpenStudy (anonymous):

I'm pretty sure 2,3,5,7 are the only numbers that can become p,q, or r, because once one of the variables become 11, the number becomes too big.

OpenStudy (anonymous):

Since neither one of us has any idea on this, could we instead work on Let a and b be real numbers such that 2^(a+6)-2^(a-6) / 2^(b+3)+ 2^(b-3) = 2016. Find a and b.

ganeshie8 (ganeshie8):

\[\dfrac{2^{a+6}-2^{a-6}}{2^{b+3}+2^{b-3}}=2016\\~\\~\\\dfrac{2^a(2^6+2^{-6})}{2^b(2^3+2^{-3})} = 2016\\~\\~\\ 2^{a-b} = \dfrac{2016(2^3+2^{-3})}{2^6+2^{-6}} \]

ganeshie8 (ganeshie8):

next you may take log both sides to isolate \(a-b\)